5 IS TO 12 AS X IS TO 48. WHAT DOES X EQUAL?
Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

48 is four times 12, so X is four times 5

Answer by stanbon(75887)

5 IS TO 12 AS X IS TO 48. WHAT DOES X EQUAL?
----------
5/12 = x/48
x = 48(5/12)
x = 4*5
x = 20
==========
Cheers,
Stan H.